Joe could not remember the product of 6 x 7, so he used a strategy to find the answer. First, he knew that 6 x 2 = 12 and then he added this product to 6 times what other number? 

Show your work below




1
Expert's answer
2021-03-10T11:30:31-0500

6*2+6*x=42

6*x=42-12

6*x=30

x=30/6=5

Other number is 5.